# On the Sharp Input-Output Analysis of Nonlinear Systems under Adversarial Attacks

### v1

@@ 4 Numerical Experiments
In this section, we provide the numerical experiments that show the effectiveness of the $\ell_{1}$-norm estimator and illustrate how the results align with our theoretical findings. To this end, we consider the following dynamics with the states $x_{t}\in\mathbb{R}^{100}$, the inputs $u_{t}\in\mathbb{R}^{5}$, the disturbances $w_{t}\in\mathbb{R}^{100}$, and the observations $y_{t}\in\mathbb{R}^{10}$ for $t=0,\dots,T-1$:
@@
(a) Comparison of Least-squares and $\ell_{1}$-norm estimator
(b) Analysis of $\tau$ and $\rho$ effects under Gaussian inputs
@@
Experiment 1. The first experiment is to compare the $\ell_{1}$-norm estimator with the commonly used least-squares estimator. The attack probability is set to $\frac{1}{2\tau+1}$, with Gaussian attack $w_{t}$ designed to have a covariance $25I_{100}$ and a mean vector whose entries are either $300$ or $1000$ depending on the sign of the corresponding coordinate of $x_{t}$. In Figure 1(a), we show that the least-squares method is vulnerable to attacks and fails to recover the system, while the $\ell_{1}$-norm estimator closely identifies the system after finite time. We provide the results under both Gaussian inputs $N(0,100I_{5})$ and non-Gaussian inputs $\text{Unif}[-8,10]^{5}$. Our theory only requires Assumptions 2.3 and 2.6 on the inputs, which is supported by our results showing that the $\ell_{1}$-norm estimator converges to the same stable region for the Gaussian inputs even when using the nonzero-mean non-Gaussian inputs (see Table 1).
Experiment 2. We now provide the experiments under different hyperparameters: the contraction factor $\rho$ and the input memory length $\tau$. Figures 1(b) and  1(c) respectively demonstrate how the estimation error evolves over time under Gaussian and non-Gaussian inputs considered in Experiment 1. The figures illustrate that a larger $\rho$ results in a higher estimation error, while a larger $\tau$ leads to a smaller eventual estimation error. These two observations align precisely with Theorem 3.1, which states an estimation error of $\Theta(\rho^{\tau})$—increasing with $\rho$ and decreasing with $\tau$. It is worth noting that this estimation error does not decay over time in the figures, which strongly supports the lower bound $\Theta(\rho^{\tau})$ established in Theorem 3.4. In Appendix F, we provide experimental details along with additional results for the case where an unbounded function is designed as the activation function $\sigma$.
@@ 5 Conclusion
In this paper, we study the identification of the input-output mappings of nonlinear dynamical systems, where control inputs are not necessarily Gaussian and the disturbances are potentially adversarial. We formulate a time-invariant input-output mapping using a linear combination of basis functions taking the input history, where we decouple the control inputs and disturbances. We propose a problem class that accurately identifies the input-output mapping, characterized by a restriction on the attack probability. We then prove that the estimation error using $\ell_{1}$-norm estimator amounts to $\Theta(\rho^{\tau})$ under the presence of adversarial attacks and show that this bound is optimal by providing a matching lower bound. Future directions include extending our analysis to a nonparametric approach under the same assumptions, where the estimator inherently involves an infinite-dimensional problem such as optimization over a function class.
Acknowledgments and Disclosure of Funding
@@ Appendix A Details on Related Works
Fully and Partially Observed Systems. In system identification, based on the degree of state observability, systems are often categorized as fully observed and partially observed systems. In fully observed systems, all states are measured, thus the outputs are identical to the states. In such systems, numerous methods have been proposed to recover the underlying system, e.g., least-squared methods [Simchowitz et al., 2018, Faradonbeh et al., 2018, Jedra and Proutiere, 2020], $\ell_{2}$- norm estimator [Yalcin et al., 2024, Zhang et al., 2024], and $\ell_{1}$-norm estimator [Kim and Lavaei, 2024]. However, in many real-world applications—such as robotics [Lauri et al., 2023], healthcare [Alagoz, 2014], and safety-critical systems [Bensoussan, 1992]—not all states are observable, giving rise to the partially observed system setting. In this case, system identification becomes substantially more challenging. A growing body of research has addressed this challenge: for instance, Sarkar et al. [2021], Oymak and Ozay [2022] identified the systems using least-squares methods, assuming disturbances are independent and follow Gaussian or sub-Gaussian distribution with zero-mean. Simchowitz et al. [2019] extended the least-squares method to setups where the disturbances can be predicted based on past observations. However, little research has been conducted in the partially observed systems when the disturbances are adversarially selected. Only recently, Kim and Lavaei [2025] investigated system identification using $\ell_{1}$-norm estimator and allowed adversarial disturbances.
@@
Nonparametric and Parametric approaches. Nonparametric approaches operate over infinite-dimensional function spaces, often leveraging techniques such as kernel methods and deep learning to model complex system dynamics [Greblicki and Pawlak, 2008, Ziemann et al., 2022]. These methods are highly flexible, making them well-suited for capturing behaviors without strong structural assumptions. However, they often come with significant computational overhead and reduced interpretability. In contrast, parametric methods approximate the system using a finite set of basis functions, typically chosen based on prior knowledge or structural insights. This approach yields models that are more interpretable, computationally efficient, and easier to analyze—especially when the chosen function class aligns well with the true system dynamics [Chen, 1995, Giannakis and Serpedin, 2001]. Moreover, the parametric framework facilitates model selection and regularization, enabling effective control over model complexity and reducing the risk of overfitting through techniques such as cross-validation or penalization.
@@ Appendix B Preliminaries on sub-Gaussian variables
In this work, we consider both inputs and attacks on the system to be sub-Gaussian variables in which the tail event rarely occurs. We use the definition given in Vershynin [2018].
@@
$\mathbb{E}\left[|w|^{k}\right]\leq(C_{1}\sqrt{k})^{k}\quad\forall k=1,2,\dots,$ (26)
$\mathbb{P}(|w|\geq s)\leq 2\exp(-s^{2}/C_{2}^{2}),\quad\forall s\geq 0,$ (27)
$\mathbb{E}[\exp(\theta w)]\leq\exp(\theta^{2}C_{3}^{2}),\quad\forall\theta\in\mathbb{R}\quad\text{if }\mathbb{E}[w]=0,$ (28)
where $C_{1},C_{2},C_{3},$ and $\|w\|_{\psi_{2}}$ are positive absolute constants that differ from each other by at most an absolute constant factor. For example, there exist $K,\tilde{K}>0$ such that $c_{1},c_{2},c_{3}\leq K\|w\|_{\psi_{2}}$ and $\|w\|_{\psi_{2}}\leq\tilde{K}\max\{c_{1},c_{2},c_{3}\}$. Note that the property (27) is also called Hoeffding’s inequality, which can be split into two inequalities if $\mathbb{E}[w]=0$:
@@
We introduce the following useful lemmas to analyze the sum of independent noncentral sub-Gaussians [Vershynin, 2018].
@@
Let $w_{1},\dots,w_{N}$ be independent, mean zero, sub-Gaussian random variables. Then, $\sum_{i=1}^{N}w_{i}$ is also sub-Gaussian and its sub-Gaussian norm is $\Theta\bigr{(}(\sum_{i=1}^{N}\|w_{i}\|_{\psi_{2}}^{2})^{1/2}\bigr{)}$.
@@
Let $\bm{W}_{t}^{(\tau,i)}$ denote the $i^{th}$ coordinate of $\bm{W}_{t}^{(\tau)}$ for $i=1,\dots,r$. Suppose that $\bm{x}_{t}^{(\tau)}=0$ for all $t$. Then, $G^{*}$ is the unique solution to the $\ell_{1}$-norm estimator (12) if
$\sum_{t=\tau}^{T-1}\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau,i)}=0\}+s^{T}\Phi(\bm{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})>0,\quad\forall s\in\mathbb{S}^{M-1}$ (34)
holds for all $i\in\{1,\dots,r\}$.
@@
Since $\bm{x}_{t}^{(\tau)}=0$, an equivalent condition for $G^{*}$ to be the unique solution of the convex optimization problem (12) is the existence of some $\epsilon>0$ such that
$\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}\|_{1}<\sum_{t=\tau}^{T-1}\|\Delta\cdot\Phi(\mathbf{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{1},\quad\forall\Delta\in\mathbb{R}^{r\times M}:0<\|\Delta\|_{F}\leq\epsilon,$ (35)
since a strict local minimum in convex problems implies the unique global minimum. A sufficient condition for (35) is to satisfy all coordinate-wise inequalities. That is, if there exist $\epsilon_{1},\dots,\epsilon_{r}>0$ such that
$\sum_{t=\tau}^{T-1}|\bm{W}_{t}^{(\tau,i)}|<\sum_{t=\tau}^{T-1}|\Delta_{i}^{T}\Phi(\mathbf{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau,i)}|,\quad\forall\Delta_{i}\in\mathbb{R}^{2km}:0<\|\Delta_{i}\|_{2}\leq\epsilon_{i}$ (36)
for all $i\in\{1,\dots,r\}$, then the inequality (35) is satisfied. For every $i$, consider a sufficiently small $\epsilon_{i}>0$. Then, for $\bm{W}_{t}^{(\tau,i)}\neq 0$, we have
$|\Delta_{i}^{T}\Phi(\mathbf{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau,i)}|$ $=(\Delta_{i}^{T}\Phi(\mathbf{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau,i)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})$
$=\Delta_{i}^{T}\Phi(\mathbf{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})+|\bm{W}_{t}^{(\tau,i)}|.$ (37)
Substituting (D) into (36) yields
$\sum_{t=\tau}^{T-1}|\bm{W}_{t}^{(\tau,i)}|<\sum_{t=\tau}^{T-1}|\Delta_{i}^{T}\Phi(\mathbf{U}_{t}^{(\tau)})|\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau,i)}=0\}+\Delta_{i}^{T}\Phi(\mathbf{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})+|\bm{W}_{t}^{(\tau,i)}|,$
which can be simplified to
$0<\sum_{t=\tau}^{T-1}|\Delta_{i}^{T}\Phi(\mathbf{U}_{t}^{(\tau)})|\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau,i)}=0\}+\Delta_{i}^{T}\Phi(\mathbf{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)}),$ (38)
for all $0<\|\Delta_{i}\|_{2}\leq\epsilon_{i}$. For every $i$, dividing both sides by $\|\Delta_{i}\|_{2}>0$ leads to the set of inequalities in (34). ∎
For the following lemma, we define the set of attack-free time as
$\mathcal{N}_{T}=\left\{t\in[\tau,T-1]:\bm{W}_{t}^{(\tau)}\equiv 0\right\}.$ (39)
We also define the set
$\mathcal{\tilde{N}}_{T}=\mathcal{N}_{T}~{}\cap~{}\Bigr{\{}(\tau+1)k:k=1,2,\dots,\Bigr{\lfloor}\frac{T-1}{\tau+1}\Bigr{\rfloor}\},$ (40)
which is the set of consecutive attack-free time at every $(\tau+1)^{\text{th}}$ timestep.
@@
Suppose that Assumptions 2.3 and 2.6 hold. Define $\nu:=\frac{\sqrt{M}\tau L_{\phi}\sigma_{u}}{\lambda}$. Then, for a fixed $s\in\mathbb{S}^{M-1}$, we have
$\mathbb{P}\biggr{(}\sum_{t=\tau}^{T-1}\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau,i)}=0\}\geq\tilde{N}_{T}\cdot\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}\biggr{)}\geq 1-\exp(-\Theta(\tilde{N}_{T})),$ (41)
where $\tilde{N}_{T}$ is the cardinality of $\mathcal{\tilde{N}}_{T}$.
@@
We first analyze the sub-Gaussian norm of $|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|$. From (8) in Assumption 2.6 with $\Phi(0)=0$, we have
$|\phi_{i}(\bm{U}_{t}^{(\tau)})|=|\phi_{i}(\bm{U}_{t}^{(\tau)})-\phi_{i}(0)|\leq L_{\phi}(\|u_{t}\|_{2}+\cdots+\|u_{t-\tau}\|_{2}).$ (42)
Due to the definition (25) of $\psi_{2}$-norm, it follows that
$\left\||\phi_{i}(\bm{U}_{t}^{(\tau)})|\right\|_{\psi_{2}}\leq\left\|L_{\phi}(\|u_{t}\|_{2}+\cdots+\|u_{t-\tau}\|_{2})\right\|_{\psi_{2}}.$
We also have
$\left\|\|u_{t}\|_{2}+\dots+\|u_{t-\tau}\right\|_{2}\|_{\psi_{2}}\leq\|\|u_{t}\|_{2}\|_{\psi_{2}}+\dots+\|\|u_{t-\tau}\|_{2}\|_{\psi_{2}}=\tau\sigma_{u}$ (43)
due to the triangle inequality. Thus, one can obtain
$\left\||s^{T}\Phi(\bm{U}_{t}^{(\tau)})|\right\|_{\psi_{2}}$ $\leq\left\|\sum_{i=1}^{M}|s_{i}\phi_{i}(\bm{U}_{t}^{(\tau)})|\right\|_{\psi_{2}}\leq\sum_{i=1}^{M}\left\||s_{i}\phi_{i}(\bm{U}_{t}^{(\tau)})|\right\|_{\psi_{2}}$
$=\sum_{i=1}^{M}|s_{i}|\left\||\phi_{i}(\bm{U}_{t}^{(\tau)})|\right\|_{\psi_{2}}\leq\sqrt{M}L_{\phi}\cdot\tau\sigma_{u}$ (44)
due to the triangle inequality and the Cauchy-Schwarz inequality.
Then, due to the property (26), we have $\mathbb{E}\left[|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|^{3}\right]\leq\Theta((\sqrt{M}\tau L_{\phi}\sigma_{u})^{3})$.
@@
$\mathbb{E}\left[|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|^{2}\right]=\mathbb{E}[s^{T}\Phi(\bm{U}_{t}^{(\tau)})\Phi(\bm{U}_{t}^{(\tau)})^{T}s]=s^{T}\mathbb{E}[\Phi(\bm{U}_{t}^{(\tau)})\Phi(\bm{U}_{t}^{(\tau)})^{T}]s\geq\lambda^{2}.$
@@
$\mathbb{E}\left[|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|^{2}\right]^{2}\leq\mathbb{E}[|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|]\cdot\mathbb{E}\left[|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|^{3}\right]$ (45)

### v2

@@ 4 Numerical Experiments
In this section, we provide the numerical experiments that show the effectiveness of the $\ell_{2}$-norm estimator and illustrate how the results align with our theoretical findings. To this end, we consider the following dynamics with the states $x_{t}\in\mathbb{R}^{100}$, the inputs $u_{t}\in\mathbb{R}^{5}$, the disturbances $w_{t}\in\mathbb{R}^{100}$, and the observations $y_{t}\in\mathbb{R}^{10}$ for $t=0,\dots,T-1$:
@@
(a) Comparison of Least-squares and $\ell_{2}$-norm estimator
(b) Comparison of the $\ell_{\alpha}$-norm estimators ($\alpha=1,2,\infty$)
@@
Experiment 1. The first experiment is to compare the $\ell_{2}$-norm estimator with the commonly used least-squares estimator. The attack probability is set to $\frac{1}{2\tau+1}$, with sub-Gaussian attack $w_{t}$ designed to have a covariance $25I_{100}$ and a mean vector whose entries are either $300$ or $1000$ depending on the sign of the corresponding coordinate of $x_{t}$. In Figure 2(a), we show that the least-squares method is vulnerable to attacks and fails to recover the system, while the $\ell_{2}$-norm estimator closely identifies the system after finite time. We provide the results under both Gaussian inputs $N(0,100I_{5})$ and non-Gaussian (nonzero-mean) inputs $\text{Unif}[-8,10]^{5}$. Our theory only requires Assumptions 2.4 and 2.5 on the inputs, which is supported by our results showing that the $\ell_{2}$-norm estimator converges to the same stable region for the Gaussian inputs even when using the nonzero-mean non-Gaussian inputs (see Table 1).
Experiment 2. Under the same experimental settings, we now present experiments comparing the $\ell_{\alpha}$-norm estimators, where $\alpha=1,2,\infty$. As discussed in Remark 3.4, all norm estimators are expected to recover the true matrix $G^{*}$ to some extent, but only the $\ell_{2}$-norm estimator theoretically achieves the optimal error $O(\rho^{\tau})$ that matches the lower bound $\Omega(\rho^{\tau})$ (see Theorems 3.1 and 3.5). Figure 2(b) indeed verifies that the $\ell_{2}$-norm estimator outperforms the other norm estimators, although the empirical differences are relatively small.
Experiment 3. We finally provide experiments under different hyperparameters: the contraction factor $\rho$ and the input memory length $\tau$, using the $\ell_{2}$-norm estimator. Figure 2(c) demonstrates how the estimation error evolves over time under non-Gaussian inputs considered in Experiment 1. The figure illustrates that a larger $\rho$ results in a higher estimation error, while a larger $\tau$ leads to a smaller eventual estimation error. These two observations align precisely with an estimation error of $O(\rho^{\tau})$—increasing with $\rho$ and decreasing with $\tau$. It is worth noting that this estimation error does not decay over time in the figures, which strongly supports the constant lower bound $\Omega(\rho^{\tau})$. In Appendix F, we provide experimental details along with additional results for the case where an unbounded function is designed as the activation function $\sigma$.
@@ 5 Conclusion
In this paper, we study the identification of the input-output mappings of nonlinear dynamical systems, where control inputs are not necessarily Gaussian and the disturbances are potentially adversarial. We formulate a time-invariant input-output mapping using a linear combination of basis functions taking the input history, where we decouple the control inputs and disturbances. We propose a problem class that accurately identifies the input-output mapping, characterized by a restriction on the attack probability. We then prove that the estimation error using $\ell_{2}$-norm estimator amounts to $O(\rho^{\tau})$ under the presence of adversarial attacks and show that this bound is optimal by providing a matching lower bound $\Omega(\rho^{\tau})$. Future directions include extending our analysis to a nonparametric approach under the same assumptions, where the estimator inherently involves an infinite-dimensional problem such as optimization over a function class.
Acknowledgments
@@ Appendix A Details on Related Works
Fully and Partially Observed Systems. In system identification, based on the degree of state observability, systems are often categorized as fully observed and partially observed systems. In fully observed systems, all states are measured, thus the outputs are identical to the states. In such systems, numerous methods have been proposed to recover the underlying system, e.g., least-squared methods (Simchowitz et al., 2018; Faradonbeh et al., 2018; Jedra & Proutiere, 2020), $\ell_{2}$-norm estimator (Yalcin et al., 2024; Zhang et al., 2025), and $\ell_{1}$-norm estimator (Kim & Lavaei, 2025a). However, in many real-world applications—such as robotics (Lauri et al., 2023), healthcare (Alagoz, 2014), and safety-critical systems (Bensoussan, 1992)—not all states are observable, giving rise to the partially observed system setting. In this case, system identification becomes substantially more challenging. A growing body of research has addressed this challenge: for instance, Sarkar et al. (2021); Oymak & Ozay (2022) identified the systems using least-squares methods, assuming disturbances are independent and follow Gaussian or sub-Gaussian distribution with zero-mean. Simchowitz et al. (2019) extended the least-squares method to setups where the disturbances can be predicted based on past observations. However, little research has been conducted in the partially observed systems when the disturbances are adversarially selected. Only recently, Kim & Lavaei (2025b; c) investigated system identification using the $\ell_{2}$-norm or $\ell_{1}$-norm estimator and allowed adversarial disturbances.
@@
Nonparametric and Parametric approaches. Nonlinear system identification approaches can generally be classified into two broad categories: nonparametric and parametric methods. Nonparametric approaches operate over infinite-dimensional function spaces, often leveraging techniques such as kernel methods and deep learning to model complex system dynamics (Greblicki & Pawlak, 2008; Ziemann et al., 2022). These methods are highly flexible, making them well-suited for capturing behaviors without strong structural assumptions. However, they often come with significant computational overhead and reduced interpretability. In contrast, our approach is based on parametric methods that approximate the system using a finite set of basis functions, typically chosen based on prior knowledge or structural insights. This approach yields models that are more interpretable, computationally efficient, and easier to analyze—especially when the chosen function class aligns well with the true system dynamics (Chen, 1995; Giannakis & Serpedin, 2001). Moreover, the parametric framework facilitates model selection and regularization, enabling effective control over model complexity and reducing the risk of overfitting through techniques such as cross-validation or penalization.
Finite-memory approximation. For a tractable identification of input-output mappings, we adopt a finite-memory approximation strategy with length $\tau$, consistent with classical system identification techniques such as Volterra series truncations (Boyd & Chua, 1985) and NARMAX models (Billings, 2013). These methods are grounded in the assumption that the dynamics of a nonlinear system can effectively be represented using a fixed window of past inputs.
Input-output mapping. In many cases, it suffices to focus on the input-output relationship—how control actions affect observable outcomes—rather than attempting to recover the full latent state dynamics (Abbeel et al., 2006; Deisenroth & Rasmussen, 2011). Similarly, to identify the input-output mapping, we design the basis functions to depend solely on the control inputs, thereby decoupling the control inputs and disturbances. This separation has proven effective and is widely adopted in various settings. For example, in model-based reinforcement learning (RL), it is common to alternate between system identification and control policy design, where the agent first learns an input-output model of the environment and then uses it to make informed decisions (Moerland et al., 2023). This simplification is particularly valuable in high-stakes applications like autonomous driving, where control inputs such as throttle and steering are mapped to observations such as heading direction, position, and velocity (Paden et al., 2016).
@@ Appendix B Preliminaries on sub-Gaussian variables
In this work, we consider both inputs and attacks on the system to be sub-Gaussian variables in which the tail event rarely occurs. We use the definition given in Vershynin (2018).
@@
$\mathbb{E}\left[|w|^{k}\right]\leq(C_{1}\sqrt{k})^{k}\quad\forall k=1,2,\dots,$ (24a)
$\mathbb{P}(|w|\geq s)\leq 2\exp(-s^{2}/C_{2}^{2}),\quad\forall s\geq 0,$ (24b)
$\mathbb{E}[\exp(\theta w)]\leq\exp(\theta^{2}C_{3}^{2}),\quad\forall\theta\in\mathbb{R}\quad\text{if }\mathbb{E}[w]=0,$ (24c)
where $C_{1},C_{2},C_{3},$ and $\|w\|_{\psi_{2}}$ are positive absolute constants that differ from each other by at most an absolute constant factor. For example, there exist $K,\tilde{K}>0$ such that $c_{1},c_{2},c_{3}\leq K\|w\|_{\psi_{2}}$ and $\|w\|_{\psi_{2}}\leq\tilde{K}\max\{c_{1},c_{2},c_{3}\}$. Note that the property (24b) is also called Hoeffding’s inequality, which can be split into two inequalities if $\mathbb{E}[w]=0$:
@@
We introduce the following useful lemmas to analyze the sum of independent noncentral sub-Gaussians (Vershynin, 2018).
@@
Let $w_{1},\dots,w_{N}$ be independent, mean zero, sub-Gaussian random variables. Then, $\sum_{i=1}^{N}w_{i}$ is also sub-Gaussian and its sub-Gaussian norm is $O\bigr((\sum_{i=1}^{N}\|w_{i}\|_{\psi_{2}}^{2})^{1/2}\bigr)$.
@@
We finally define a notion of subexponential, which is essentially a squared sub-Gaussian.
Lemma B.5.
$w$ is sub-Gaussian if and only if $w^{2}$ is subexponential, and it holds that
$\|w\|_{\psi_{2}}^{2}=\|w^{2}\|_{\psi_{1}},$
where the $\psi_{1}$-norm is defined as
$\|w^{2}\|_{\psi_{1}}=\inf\left\{c>0:\mathbb{E}\left[\exp\left(\frac{w^{2}}{c}\right)\right]\leq 2\right\}.$ (28)
@@
For convenience, we define $\mathbb{I}_{\pm}(\cdot)$ as the indicator function that equals $1$ if the event occurs and 1-1 otherwise.
@@
Suppose that $\bm{x}_{t}^{(\tau)}=0$ and $\bm{\epsilon}_{t}=0$ for all $t$. Then, $G^{*}$ is the unique solution to the $\ell_{2}$-norm estimator (12) if
$\sum_{t=\tau}^{T-1}\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}>0$ (31)
holds for all $Z\in\mathbb{R}^{r\times M}$ such that $\|Z\|_{F}=1$.
@@
Since $\bm{x}_{t}^{(\tau)}=0$ and $\bm{\epsilon}_{t}=0$, an equivalent condition for $G^{*}$ to be the unique solution of the convex optimization problem (12) is the existence of some $\bar{\Delta}>0$ such that
$\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}\|_{2}<\sum_{t=\tau}^{T-1}\|\Delta\cdot\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{2},\quad\forall\Delta\in\mathbb{R}^{r\times M}:0<\|\Delta\|_{F}\leq\bar{\Delta},$ (32)
since a strict local minimum in convex problems implies the unique global minimum. Observe that we have
$\sum_{t=\tau}^{T-1}\|\Delta\cdot\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{2}-\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}\|_{2}$
$\hskip 2.84526pt=\sum_{t=\tau}^{T-1}\mathbb{I}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\|\Delta\cdot\Phi(\bm{U}_{t}^{(\tau)})\|_{2}+\mathbb{I}\{\bm{W}_{t}^{(\tau)}\neq 0\}\cdot\left(\|\Delta\cdot\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{2}-\|\bm{W}_{t}^{(\tau)}\|_{2}\right)$
$\hskip 2.84526pt\geq\sum_{t=\tau}^{T-1}\mathbb{I}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\|\Delta\cdot\Phi(\bm{U}_{t}^{(\tau)})\|_{2}+\mathbb{I}\{\bm{W}_{t}^{(\tau)}\neq 0\}\cdot\left(-\|\Delta\cdot\Phi(\bm{U}_{t}^{(\tau)})\|_{2}\right)$
$\hskip 2.84526pt=\sum_{t=\tau}^{T-1}\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|\Delta\cdot\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}.$ (33)
Thus, a sufficient condition for (32) is to satisfy
$\sum_{t=\tau}^{T-1}\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|\Delta\cdot\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}>0,\quad\forall\Delta\in\mathbb{R}^{r\times M}:0<\|\Delta\|_{F}\leq\bar{\Delta}.$ (34)
For each $\Delta$, dividing both sides of (34) by $\|\Delta\|_{F}>0$ leads to the set of inequalities in (31). ∎
We will now analyze the sub-Gaussian norm of $\|\bm{U}_{t}^{(\tau)}\|_{2}$.
@@
Under Assumption 2.5, we have $\bigr\|\|\bm{U}_{t}^{(\tau)}\|_{2}\bigr\|_{\psi_{2}}\leq\sqrt{\tau+1}\cdot\sigma_{u}$.
@@
We will use Lemma B.5, which connects sub-Gaussian and subexponential variables. Since the sub-Gaussian norms of $\|u_{t}\|_{2}$ for all $t$ are bounded by $\sigma_{u}$, we equivalently have
$\left\|\|u_{t}\|_{2}^{2}\right\|_{\psi_{1}}\leq\sigma_{u}^{2},\quad\forall t\geq 0.$
It follows that
$\left\|\|\bm{U}_{t}^{(\tau)}\|_{2}^{2}\right\|_{\psi_{1}}=\left\|\sum_{i=0}^{\tau}\|u_{t-i}\|_{2}^{2}\right\|_{\psi_{1}}\leq\sum_{i=0}^{\tau}\left\|\|u_{t-i}\|_{2}^{2}\right\|_{\psi_{1}}\leq(\tau+1)\sigma_{u}^{2}.$
We again hinge on Lemma B.5 to arrive at the conclusion. ∎
Lemma D.3.
Suppose that Assumptions 2.4, 2.5, and 2.8 hold. Define $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$. Then, for a fixed $Z\in\mathbb{R}^{r\times M}$ such that $\|Z\|_{F}=1$, we have
$\mathbb{E}\left[\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}\right]=\Omega\left(\frac{(2(1-p)^{\tau}-1)\cdot\lambda}{\nu^{3}}\right).$ (35)
Proof.
We first analyze the sub-Gaussian norm of $\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}$. From (9) in Assumption 2.4 with $\Phi(0)=0$, we have
$|\phi_{i}(\bm{U}_{t}^{(\tau)})|=|\phi_{i}(\bm{U}_{t}^{(\tau)})-\phi_{i}(0)|\leq L_{\phi}\|\bm{U}_{t}^{(\tau)}\|_{2}.$ (36)
Due to Lemma D.2, it follows that
$\left\||\phi_{i}(\bm{U}_{t}^{(\tau)})|\right\|_{\psi_{2}}\leq L_{\phi}\left\|\|\bm{U}_{t}^{(\tau)}\|_{2}\right\|_{\psi_{2}}=L_{\phi}\sqrt{\tau+1}\cdot\sigma_{u}$
Thus, one can obtain
$\left\|\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}\right\|_{\psi_{2}}$ $\leq\left\|\sum_{i=1}^{M}\|z_{i}\phi_{i}(\bm{U}_{t}^{(\tau)})\|_{2}\right\|_{\psi_{2}}\leq\sum_{i=1}^{M}\left\|\|z_{i}\phi_{i}(\bm{U}_{t}^{(\tau)})\|_{2}\right\|_{\psi_{2}}$
$=\sum_{i=1}^{M}\|z_{i}\|_{2}\left\||\phi_{i}(\bm{U}_{t}^{(\tau)})|\right\|_{\psi_{2}}\leq\sqrt{M}L_{\phi}\sqrt{\tau+1}\cdot\sigma_{u},$ (37)
where $z_{i}$ is the $i^{\text{th}}$ column of $Z$. The inequalities follow from the triangle inequality and the Cauchy-Schwarz inequality.
Then, due to the property (24a), we have $\mathbb{E}\left[\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{3}\right]=O((\sqrt{M\tau}L_{\phi}\sigma_{u})^{3})$.
@@
$\mathbb{E}\left[\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{2}\right]$ $=\mathbb{E}[\text{trace}(Z^{T}Z\cdot\Phi(\bm{U}_{t}^{(\tau)})\Phi(\bm{U}_{t}^{(\tau)})^{T})]$
$=\text{trace}(Z^{T}Z\cdot\mathbb{E}[\Phi(\bm{U}_{t}^{(\tau)})\Phi(\bm{U}_{t}^{(\tau)})^{T}])\geq\lambda^{2}\cdot\text{trace}(Z^{T}Z)=\lambda^{2}.$ (38)
@@
$\mathbb{E}\left[\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{2}\right]^{2}\leq\mathbb{E}\left[\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}\right]\cdot\mathbb{E}\left[\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{3}\right]$ (39)